OpenStudy (anonymous):

Find the value of the variable if P is between J & K. JP = 3y+1 ; PK = 12y - 4 ; JK = 75 Choose one answer. a. 5.2 b. 8.7 c. 3.9 d. 4.3

OpenStudy (anonymous):

Note that JP + PK = JK so, 3y + 1 + 12y - 4 = 75 15y - 3 = 75 15y = 78 y = 78/15 or 5.2

OpenStudy (anonymous):

Thank you!!